OpenStudy (anonymous):

At t= 0, a particle starts from rest at x= 0, y = 0, and moves in the xy plane with an acceleration a⃗ = (4.0i^ + 3.0j^)m/s2. Assume t is in seconds. Determine the x component of velocity as a function of time t. etermine the ycomponent of velocity as a function of time t.

OpenStudy (anonymous):

\[\bar a=d \bar v/dt \rightarrow \bar v=\int\limits\limits_{}^{}\bar a dt=\int\limits\limits_{}^{}(4.0 \hat i+3.0 \hat j)dt+\bar C=(4t+C_x) \hat i+(3t+C_y)\hat j\]Vx=4t+Cx Vy=3t+Cy As at t=0 Vx=0 and Vy=0 then Cx=0 and Cy=0 and we have: Vx=4t Vy=3t